# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

### v1

@@ 6.2. Phase transition of $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $M=3N+3$
However, in either case the component sequence $d_{m+1}^{1}d_{m+2}^{1}\ldots$ ends with $((N+2)N(N+1))^{\infty}={\mathbf{v}}_{1}^{\infty}$, which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. Similarly, if $d_{m+1}$ equals any other digit in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+2)(N+1)}),\theta^{j}(\alpha_{(N+2)N})\right\}$, we can deduce the same contradiction. So, $d_{m+1}=\alpha_{(N+1)(N+1)}$; and therefore by the same argument as above we can prove that $d_{m+1}d_{m+2}\ldots=\alpha_{(N+1)(N+1)}^{\infty}={\mathbf{c}}_{0}^{\infty}$.
@@
which implies that $d_{m+1}^{1}d_{m+2}^{1}\ldots={\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. If $d_{m+1}d_{m+2}d_{m+3}=\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)}$, then by (6.16) and Proposition 2.3 we obtain that
@@
which implies that $d_{m+4}^{1}d_{m+5}^{1}\ldots$ ends with ${\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, again leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. So, we must have (6.17).
@@
If $d_{m+1}$ is taken some other digit from $\bigcup_{k=0}^{2}\left\{\theta^{k}(\alpha_{(N+2)(N+1)}),\theta^{k}(\alpha_{(N+2)N})\right\}$, we can prove by the same argument as above that $d_{m+1}d_{m+2}\ldots$ ends with $\theta^{k}({\mathbf{c}}_{1}^{\infty})$ for some $k=0,1,2$. This completes the proof. ∎
@@
If $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k+1}^{+}$, the by Proposition 6.3 and (6.12) it follows that
@@
which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@ Acknowledgements
The authors thank Wolfgang Steiner for his insight on an alternate definition of the Thue-Morse type sequences. The first author was supported by NSFC No. 12301108. The second author was supported by Chongqing NSF: CQYC20220511052. The third author was supported by NSFC No. 12071148 and 12471085.

### v2

@@ 6.2. Phase transition of $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $M=3N+3$
However, in either case the component sequence $d_{m+1}^{1}d_{m+2}^{1}\ldots$ ends with $((N+2)N(N+1))^{\infty}={\mathbf{v}}_{1}^{\infty}$, which leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. Similarly, if $d_{m+1}$ equals any other digit in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+2)(N+1)}),\theta^{j}(\alpha_{(N+2)N})\right\}$, we can deduce the same contradiction. So, $d_{m+1}=\alpha_{(N+1)(N+1)}$; and therefore by the same argument as above we can prove that $d_{m+1}d_{m+2}\ldots=\alpha_{(N+1)(N+1)}^{\infty}={\mathbf{c}}_{0}^{\infty}$.
@@
which implies that $d_{m+1}^{1}d_{m+2}^{1}\ldots={\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. If $d_{m+1}d_{m+2}d_{m+3}=\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)}$, then by (6.16) and Proposition 2.3 we obtain that
@@
which implies that $d_{m+4}^{1}d_{m+5}^{1}\ldots$ ends with ${\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, again leading to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. So, we must have (6.17).
@@
If we choose $d_{m+1}$ for some other digit from $\bigcup_{k=0}^{2}\left\{\theta^{k}(\alpha_{(N+2)(N+1)}),\theta^{k}(\alpha_{(N+2)N})\right\}$, then by the same argument as above it follows that $d_{m+1}d_{m+2}\ldots$ ends with $\theta^{k}({\mathbf{c}}_{1}^{\infty})$ for some $k=0,1,2$. This completes the proof. ∎
@@
If $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k+1}^{+}$, then by Proposition 6.3 and (6.12) it follows that
@@
which leads to a contradiction to $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
@@ Acknowledgements
The authors thank the anonymous referee for many useful comments which greatly improve the presentation of the paper. They also thank Wolfgang Steiner for his insight on an alternate definition of the Thue-Morse type sequences. The first author was supported by the National Natural Science Foundation of China No. 12301108. The second author was supported by the Chongqing Natural Science Foundation: CQYC20220511052 and the Scientific Research Innovation Capacity Support Project for Young Faculty No. ZYGXQNISKYCXNLZCXM-P2P. The third author was supported by the National Natural Science Foundation of China No. 12071148 and 12471085..